OpenStudy (anonymous):

determine w/out graphing whether the system of equations has exactly one solution, no solutions, or infinite number of solutions 4x + 3y =8 6y = -8x +4

OpenStudy (yuki):

one of the visually easiest ways to solve a problem like this is to fix the equations into "y = mx + b" the slope intercept form would you like to do that first ?

OpenStudy (anonymous):

thank you i just did that I had bran freeze

OpenStudy (anonymous):

Once you have them in either slope intercept, or point slope form you can find the answer by inspection.

OpenStudy (anonymous):

paralell many solution

OpenStudy (anonymous):

Is that for differential equations? It's a weird questions because they are both clearly specific linear lines.

OpenStudy (anonymous):

No, it's for algebra 2 most likely.

OpenStudy (anonymous):

Also, Nancy is wrong, sorry =(

OpenStudy (anonymous):

All right, well ill be quiet then. :)

OpenStudy (yuki):

if the slopes are the same, then there are two cases. 1), they are the same line algebraically, the two lines will look exactly the same so there are infinitely many solutions since the "two lines" intersect everywhere 2), they have the same slope but different y-intercept algebraically, m is the same but b is different they are parallel and never intersect, so there is no solution. if the slopes are different, then they will intersect at some point. so there will be exactly one solution.

OpenStudy (anonymous):

I solve they are the same slope

OpenStudy (anonymous):

Indeed, but they are not the same line. Two parallel lines that are not the same line will never intersect.

OpenStudy (anonymous):

4x + 3y =8 y= -4/3x+2 6y = -8x +4 and y= -4/3x+2 yes the same slope is paraell y=4/3x +2/3 y= -4/3x+2 and

OpenStudy (anonymous):

I forgot the same no solution,if the same line than many solution polpak right